OpenStudy (anonymous):

Parametric Differentiation x= sqrt{t} , y = sqrt{t-1} t = 2 Find dy/dx and evaluate the specified value of the parameter

OpenStudy (anonymous):

alright what throws me off is idk if to just make the y equation the sqrt of t and differentiate it using the constant rule or to say (t-1)1/2 then differentiate we're only concerned with t so its just the sqrt of t right for my y equation?

OpenStudy (amistre64):

derive both parts seperately and then put the y over the x and solve with t

OpenStudy (amistre64):

x = sqrt{t} y = sqrt{t-1} dx = 1/2sqrt(2) dy = 1/2sqrt{t-1}

OpenStudy (amistre64):

sqrt(2) means sqrt(t) :)

OpenStudy (anonymous):

dont we get rid of the 1 or kinda ignore it?

OpenStudy (amistre64):

no, its a part of the function

OpenStudy (amistre64):

the innards derive to simply '1' so it doesnt affect the normal flow of the deriving

OpenStudy (anonymous):

so it'll be1/2 (t-1)-1/2(1)

OpenStudy (amistre64):

now if it had been: y = sqrt(2x+1) ; then that derives to: D(2x+2) ---------- 2sqrt(2x+1)

OpenStudy (amistre64):

\[{dy \over dx}=\frac{1/2\sqrt{t-1}}{1/2\sqrt{t}}\]

OpenStudy (anonymous):

i know how to do the bottom im not sure about the top

OpenStudy (anonymous):

because my lecturer said were not concerned about the values with no t

OpenStudy (anonymous):

but that one is different

OpenStudy (amistre64):

get this looking normal and then rationalize the denom

OpenStudy (amistre64):

\[\frac{2\sqrt{t}}{2\sqrt{t-1}}\]

OpenStudy (amistre64):

\[\frac{\sqrt{t}}{\sqrt{t-1}}*\frac{\sqrt{t-1}}{\sqrt{t-1}}\]

OpenStudy (amistre64):

it might just have to end up having a negative square root .... :)

OpenStudy (anonymous):

why'd you put the two in fornt of the sqrt{t}

OpenStudy (amistre64):

\[\frac{\sqrt{t^2 -t}}{t-1}\]

OpenStudy (amistre64):

because sqrt(t) derives to 1/2sqrt(t)

OpenStudy (amistre64):

now when t=2 we get: sqrt(2)/2

OpenStudy (amistre64):

er..... \[\sqrt2\]

OpenStudy (amistre64):

look over what I got, I gotta go to class for a few hours :) good luck

OpenStudy (anonymous):

ok thank you so much! have fun!!!